Please explain. 0/4 pts Question 5 Electrons exist in orbitals within an atom. Bohrs model described a single electron system well. But the3/6 pts Question 15 Compare the atomic radius of Argon to each of the following. Krypton atom Argon is smaller. Calcium atom0/2 pts Question 16 Which is the most exothermic? The ionization of Ca The electron affinity of P The ionization energy of Mg

0/4 pts Question 5 Electrons exist in orbitals within an atom. Bohr's model described a single electron system well. But the energy level for orbitals change when as you add a second and more electron. In a single electron system, there are (a)_orbital(s) at the n-2 level. These n-2 orbital(s) exist at (b) sub-level(s) in the atom. In a multi-electron system, (c)_orbital(s) at the n-2 level. These n-2 orbital(s) exist at (d)sub-level(s) in the atom. Bohr was never able to explain this difference between single electron and multi electron systems, which produces line splitting in many elements line spectra. Schrödinger's quantum model predicts this splitting.
3/6 pts Question 15 Compare the atomic radius of Argon to each of the following. Krypton atom Argon is smaller. Calcium atom Argon is smaller. Chloride ion (CI1-) Argon is the same size. Titanium-4 ion (Ti4+) Argon is smaller.
0/2 pts Question 16 Which is the most exothermic? The ionization of Ca The electron affinity of P The ionization energy of Mg The electron affinity of C
Answer #1

Q-5-at n=2 level, there are 4 orbitals and 2 subshells in both single and multi-electron system.

i.e,(a)=4 (b)=2 (c)=4 (d)=2

Q-15-argon is smaller than Krypton i.e, Ar=71pm and Kr=88pm

argon is smaller than Ca atom i.e, Ar=71pm and Ca=231pm

argon is smaller than Cl- i.e, Ar=71pm and Cl-=181pm

argon is larger  than Ti4+ i.e Ar=71pm and Ti4+=51pm

Q-16-electron affinity is the exothermic process while ionization energy is an endothermic process

so out of electron affinity of Cl and P

Cl =248.5KJ/Mol and P=72KJ/Mol which means electron affinity of Cl is most exothermic.